I am asked to show how to simplify the following:
3/8x -1/6y - 1/16x + x - 1/3y
The answer that I was given is 21/16x - 1/2y.
I understand how 21/16x is derived however I do not understand how 1/2y is the answer. Can yo please help?
: I am asked to show how to simplify the following:
3/8x -1/6y - 1/16x + x - 1/3y
The answer that I was given is 21/16x - 1/2y.
I understand how 21/16x is derived however I do not understand how 1/2y is the answer. Can yo please help?

Answer by baholik06(4)

You have to combine like varibles. 3/8x-1/16x+x=21/16x -1/6y-1/3y=-1/2y
So your answer is 21/16x-1/2y